OpenStudy (anonymous):

What value for n makes this equation true? 6 • 25 = (6 • 20) + (6 • n)

OpenStudy (anonymous):

N = 5. This value makes the equation true because: 6 * 25 = 150 (6 * 20) + (6 * 5) = 150